OpenStudy (anonymous):

I don't get it at all, and the way other people explained it didn't help. Someone said that you have to multiply the percentages in decimal form, and then multiply that answer by 14. So when I do that, I get the answer of 2.94. Is that the correct answer? Do I just round up? Hanna makes her own furniture polish using thirty percent vinegar and seventy percent olive oil. Write and 70% olive oil. Write and solve an equation to determine how much total solution she made if she used 14 tablespoons of olive oil. Show your work.

OpenStudy (anonymous):

uhh... yeah. That didn't help. Uhh knock knock, come in. @Loser66

OpenStudy (jdoe0001):

hmmm Auubbzz do you know what 30% means? do you know what the 70% of something is?

OpenStudy (jdoe0001):

do you even know the meaning of a percentage of something?

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

:P

OpenStudy (anonymous):

Okay you guys obviously aren't helping. Besides, I kind of already figured it out. Do you even know how to answer a question? @jdoe0001

OpenStudy (mathstudent55):

If 14 tablespoons are 70% of a total, then the total is 14/70% = 14/0.70 = 20. She used 14 tablespoons of olive oil and 6 tablespoons of vinegar to make 20 tablesopoons of polish.

OpenStudy (pixiedust1):

Oh my gosh!!! It's the guru!!! :D HI HERO!!!

OpenStudy (ranga):

Assume total solution = x tablespoons 70% of the total solution is olive oil Therefore, olive oil = x * 70/100 = 0.7x She used 14 tablespoons of olive oil. Therefore, 0.7x = 14. Solve for x. Divide both sides by 0.7 x = 14/0.7 = 20 tablespoons.